OpenStudy (anonymous):

a line passes through a point (-2,3).find an equation of this line if it has slope (3/2)

OpenStudy (anonymous):

y=mx+b m is the slope (3/2) b is the y intercept, in this case it's 3

OpenStudy (karatechopper):

1/-5

OpenStudy (anonymous):

what is x?

OpenStudy (precal):

y=mx+b 3=(3/2)(-2)+b 3=-3+b 0=b equation is y=(3/2)x+0 or y=(3/2)x